# MARVEL: Unlocking the Multi-Modal Capability of Dense Retrieval via Visual Module Plugin

### v3

@@ Abstract
This paper proposes Multi-modAl Retrieval model via Visual modulE pLugin (MARVEL), which learns an embedding space for queries and multi-modal documents to conduct retrieval. MARVEL encodes queries and multi-modal documents with a unified encoder model, which helps to alleviate the modality gap between images and texts. Specifically, we enable the image understanding ability of the well-trained dense retriever, T5-ANCE, by incorporating the visual module’s encoded image features as its inputs. To facilitate the multi-modal retrieval tasks, we build the ClueWeb22-MM dataset based on the ClueWeb22 dataset, which regards anchor texts as queries, and exacts the related text and image documents from anchor-linked web pages. Our experiments show that MARVEL significantly outperforms the state-of-the-art methods on the multi-modal retrieval dataset WebQA and ClueWeb22-MM. MARVEL provides an opportunity to broaden the advantages of text retrieval to the multi-model scenario. Besides, we also illustrate that the language model has the ability to extract image semantics and partly map the image features to the input word embedding space. All codes are available at https://github.com/OpenMatch/MARVEL.
@@
[…]
@@ 1 Introduction
Multi-modal retrieval Bain et al. (2021); Awad et al. (2021); Arni et al. (2008); Chang et al. (2022) aims to return fusion results of images and texts to answer user questions. The task can be modeled using a divide-and-conquer pipeline Chang et al. (2022); Liu et al. (2023b) or universal dense retrieval Liu et al. (2023b). UniVL-DR Liu et al. (2023b) encodes queries and multi-modal documents into a universal embedding space for multi-model retrieval. However, this work encodes image features and texts using different encoders from CLIP Radford et al. (2021) and the separated text and image encoding leads to a modality gap in representing multi-modal documents. It makes UniVL-DR design an additional image verbalization method to alleviate the modality gap and also limits the text retrieval models Karpukhin et al. (2020); Xiong et al. (2021a); Zhan et al. (2021); Li et al. (2021b); Yu et al. (2021) to excel their advantages in multi-modal scenarios.
@@ 3.1 Preliminary of Multi-Modal Retrieval
The multi-model retrieval task requires retrievers to conduct relevance modeling, cross-modal matching, and modality fusion Liu et al. (2023b). Previous work Liu et al. (2023b) maps text and image documents in an embedding space for retrieval, encodes texts and images using different encoders, and tries to bridge the modality gap using image verbalization methods. However, this limits the capability of dense retrieval models, hindering the expansion of text matching knowledge for learning representations for multi-modal documents.
@@ 3.2 Universal Multi-Modal Encoding
We show the model architecture in Figure 2. Different from previous work Liu et al. (2023b), we can universally encode query $q$ and multi-model documents $\mathcal{D}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m},d_{\text{Image}}^{1},...,d_{\text{Image}}^{n}\}$ using one encoder, T5-ANCE-CLIP:
@@ 3.2.1 Dense Retrieval with Visual Plugin
MARVEL starts from the T5-ANCE model Yu et al. (2023), which is a dense retrieval model that is well-trained using MS MARCO dataset Nguyen et al. (2016). Then we enable T5-ANCE by incorporating the visual module from the vision-language model, CLIP Radford et al. (2021), and conduct the T5-ANCE-CLIP model. We can use a universal encoder, T5-ANCE-CLIP, to encode texts, image features and image documents.
@@ 5 Evaluation Result
In this section, we first evaluate the performance of MARVEL and then conduct ablation studies. Subsequently, we explore the effectiveness of different visual and language model fusion methods and analyze the role of image features in the MARVEL. Some case studies are shown in Appendix A.6.
@@ 5.1 Overall Performance
The multi-modal retrieval performance of MARVEL and baseline models is shown in Table 2.
@@ 5.2 Ablation Study
In the comparison between MARVEL and MARVEL (w/o CLIP Pretraining), pretraining the visual understanding module shows its effectiveness by improving the performance on single/multi-modal retrieval tasks. It shows that the image-caption alignment relations provide some opportunities to adapt the visual module to the language model via pretraining. Subsequently, MARVEL also outperforms MARVEL (w/o MS MARCO Pretraining), especially on the text retrieval task. It demonstrates that MARVEL can broaden the advantage of text relevance modeling to the multi-modal retrieval task. To unify the multi-modal encoding, MARVEL follows previous work Hannan et al. (2020) uses prompt tokens to indicate the start and end positions of encoded image features (Eq. 6), aiming to distinguish the image features from text token embeddings. These image prompt tokens bring light improvements, illustrating their roles in multi-modal document representation.
@@ 5.3 Retrieval Effectiveness of Different Visual-Language Fusion Methods
MARVEL-Concat Image 37.20 81.14
@@
Table 4: Retrieval Performance of the Models using Different Visual-Language Fusion Methods. The Sum and Concat methods are similar to previous work Liu et al. (2023b), which only replace the image caption encoder with T5-ANCE. The CLIP-Sum model is the CLIP-DPR model from previous work Liu et al. (2023b). All models are trained with in-batch negatives. MRR@10 is used to evaluate the retrieval effectiveness of all models.
@@
Different from Liu et al. (2023b), we use T5-ANCE and CLIP as the text and image encoders, respectively. These models have different architectures and are pretrained on text retrieval and image-caption matching tasks. The multi-modal retrieval performance of CLIP-Sum decreases when we encode the image caption with a stronger retrieval model T5-ANCE (MARVEL-Sum) instead of CLIP. It demonstrates that incorporating an additional visual module into a well-trained dense retrieval model is still challenging for multi-modal retrieval. Notably, MARVEL provides a promising way to enable the image understanding ability of dense retrieval models by using the visual module plugin modeling method.
@@
Figure 3: Attention Distribution of MARVEL-DPR. The attention weights of image features are shown in Figure 3(a). And the attention weight entropy of image captions and features is shown in Figure 3(b).
@@
As shown in Figure 3. The attention scores are calculated by cross attentions from the decoder to the encoder module of T5. We first show the attention weight distribution of image features in Figure 3(a). When we only finetune the language model, the attention heads tend to allocate more attention weights to image features, helping to adapt the visual module in the language model. On the other hand, the image features win more attention weights when the CLIP module is finetuned. However, as shown in Figure 3(b), only finetuning the CLIP module shows a scattered attention weight mechanism than other models, which misleads the T5-ANCE to capture more important information from encoded representations of documents. All these phenomena demonstrate the necessity of the training strategies of MARVEL, which pretrain visual module for adaption and only finetune the language model for multi-modal retrieval. In addition, we show the retrieval effectiveness with different finetuning methods in Appendix A.2.
Model MRR@10 NDCG@10 Rec@100
MARVEL-DPR 55.71 52.94 88.23
w/ 1-NN Token 38.80 35.89 73.59
w/ 5-NN Tokens 42.39 39.27 75.04
w/ Random Token 37.73 35.34 71.92
MARVEL-ANCE 65.15 62.95 92.40
w/ 1-NN Token 51.37 48.27 80.47
w/ 5-NN Tokens 52.22 49.35 81.88
w/ Random Token 44.22 41.55 71.23
5.4 Learned Semantics of Image Features
As shown in Table 5, replacing encoded image features with $k$-NN token embeddings generally outperforms the retrieval model using randomly selected token embeddings. It demonstrates that the visual plugin module effectively maps image semantics in the input space of the language model, and the ability to keep growing with more token embeddings (5-NN). However, the retrieval performance significantly drops when employing $k$-NN token embeddings to replace the image features, compared to the MARVEL model. It demonstrates the role of encoded image features beyond the semantic representations of word embeddings. The encoded image features may act as a kind of prompt, encouraging language models to capture image semantics Merullo et al. (2023). More cases are shown in Appendix A.4.
@@ 6 Conclusion
This paper proposes Multi-modAl Retrieval via Visual modulE pLugin (MARVEL). MARVEL integrates a visual plugin module to a well-trained dense retriever and pretrains the visual module with image-caption contrastive training for adaption. Our MARVEL model achieves state-of-the-art on all benchmarks by unifying the multi-modal document encoding and alleviating the modality gap between images and texts.
@@ Appendix A Appendix
A.1 Experimental Details of MARVEL Pretraining Data
@@
To pretrain the visual module in MARVEL, we collect the image-caption pairs from the ClueWeb22 dataset. We retain the English pages, extract the content within the image tag and use the image and alt-text to construct the image-caption pair. To ensure the quality of the pretraining dataset, following LAION-400M Schuhmann et al. (2021), we use CLIP to calculate the embeddings of images and captions and compute the cosine similarity between the two embeddings. Subsequently, we discard all samples with a cosine similarity lower than 0.3. The pretraining dataset contains 1.6M image-caption pairs, and we randomly select 10,000 pieces of data as the development set and use the rest for pretraining visual module.
Finetune Modaltiy MRR@10 NDCG@10 Rec@100
Text 53.99 48.36 85.70
CLIP & T5 Image 60.91 62.22 93.71
Multi 55.41 53.02 88.48
Text 54.28 48.24 85.80
T5 Image 60.81 61.74 93.55
Multi 55.58 52.86 88.50
Text 49.39 42.04 74.90
CLIP Image 55.18 55.16 87.91
Multi 44.54 41.03 75.91
@@
N/A Image 51.60 51.63 83.49
Multi 44.09 40.65 74.16
Table 6: The Retrieval Performance with Different Training Strategies. We freeze each module of MARVEL-DPR to explore the benefits of training between different modules.
A.2 Retrieval Effectiveness of Different Finetuning Strategies
@@
As shown in Table 6, finetuning the CLIP module indeed improves the retrieval performance of the whole frozen model, especially in the image retrieval task. This observation shows that multi-modal training signals are effective to benefit the capability of visual modules. When we only tune the parameters of T5, MARVEL-DPR achieves significant improvements over the frozen model, showing the language model’s strong ability to adapt the visual module to the dense retriever. Nevertheless, the fully finetuned model decreases the retrieval performance of MARVEL-DPR that only finetunes T5. It shows the necessity of the pretraining-and-then-finetuning strategy of MARVEL, which pretrains the visual understanding module for adaption and finetunes the language model for multi-modal retrieval.
A.3 More Details of ClueWeb22-MM
@@
Figure Text
“Hello”, “stocked”, “charged”, “institu”]
“Hi”, “auftrag”, “he”, “ban”, “sets”, “7,000”, “representation”]
“default”, “16”, “raum”, “haunt”, “unanimous”]
“russia”, “Color”, “vivid”, “HM”, “arbeit”, “default”]
[Uncaptioned image] Manual Caption: Green Corn Dance at Tesuque Indian Pueblo near Santa Fe, New Mexico File name: 06 10 015670
Nearest Tokens: [“parametri”, “damaged”, “members”, “Feuer”, “muddy”, “potentiel”, “membres”, “statutory”, “aires”,
“monetary”, “pun”, “Chemical”, “bookstore”, “Donna”, “tari”, “elemente”, “lawmakers”, “IN”, “bankruptcy”, “tomato”,
“chemical”, “materiale”, “banks”, “souvent”, “graphical”, “paper”, “vorgesehen”]
Nearest Tokens w/o CLIP Pretraining: [ “immediate”,“OR”, “ammunition”, “comune”, “cartoon”, “LAR”, “SO”, “LED”,
“sign”, “Sin”, “unanimous”, “entière”, “auftrag”, “historischen”, “concise”, “Arbeits”, “obligaţi”, “gray”, “vivid”, “Hi”,
“enforcement”, “part”, “UD”, “He”, “suffisamment”, “ened”, “agencies”, “EU”, “charged”, “commune”, “clown”]
@@
To show the details of our ClueWeb22-MM dataset, we show the data collection, data processing and data statistics.
@@
To further guarantee the quality of the dataset, we use T5-ANCE Yu et al. (2023) to estimate the relevance between the anchor and its corresponding image document. We encode all captions of image documents using T5-ANCE, use the anchor texts as queries to retrieve the images and reserve the anchors that are ranked in the top 10. Finally, we respectively sample 10,000 queries to build the development set and test set. The rest data are used for finetuning models, which contain 72,028 queries.
@@
A.4 Effectiveness of Image Features
We show some case studies in Table 10 to show the effectiveness of the visual understanding module by verbalizing the semantics of encoded image features using word tokens.
We randomly select five image documents of different topics and represent the encoded image feature with some tokens to verbalize the image semantics. Specifically, we first use the visual plugin modules of MARVEL and MARVEL w/o CLIP Pretraining to encode the image features. Then, to show the semantics of the encoded image features, we utilize cosine similarity to find the $k$-NN tokens for each encoded image feature. Finally, we utilize the token with the highest score to represent the semantics of the encoded image features.
Overall, compared with the MARVEL w/o CLIP Pretraining model, MARVEL learns more effective representations for image documents. These representations are closer to the word tokens associated with the semantics of the image documents. It demonstrates that MARVEL has the ability to better adapt the visual understanding module to dense retrievers.
For the first two examples, MARVEL w/o CLIP Pretraing learns more similar representations for both image documents. The related word tokens of these image documents contain lots of same tokens, such as “7,000”, “Hi”, and “RAM”, which are unrelated to the semantics of the image documents. On the contrary, our MARVEL model can learn more similar semantics to both image documents. Specifically, MARVEL verbalizes the first image document using the words “brightness”, “resident” and “store”, which are related to the image description of “Bourbon Street”. And MARVEL also learns the semantics of “animals”, “wildlife” and “creatures” of the second image document. The next two instances show the effectiveness of MARVEL in learning more fine-grained semantics of the image documents by verbalizing the image documents with more related words, such as “militari”, “vehicle”, “flag”, “legislatur”, and “government”. All these cases illustrate our visual module pretraining strategy’s important role in improving MARVEL’s visual understanding ability.
Even though our visual module pretraining method shows its effectiveness, the last case shows the limitation of existing CLIP based models. The image document describes the semantics of “Green Corn Dance”, which is still hard to comprehend by the visual understanding module.

### v4

@@ Abstract
This paper proposes Multi-modAl Retrieval model via Visual modulE pLugin (MARVEL), which learns an embedding space for queries and multi-modal documents to conduct retrieval. MARVEL encodes queries and multi-modal documents with a unified encoder model, which helps to alleviate the modality gap between images and texts. Specifically, we enable the image understanding ability of the well-trained dense retriever, T5-ANCE, by incorporating the visual module’s encoded image features as its inputs. To facilitate the multi-modal retrieval tasks, we build the ClueWeb22-MM dataset based on the ClueWeb22 dataset, which regards anchor texts as queries, and extracts the related text and image documents from anchor-linked web pages. Our experiments show that MARVEL significantly outperforms the state-of-the-art methods on the multi-modal retrieval dataset WebQA and ClueWeb22-MM. MARVEL provides an opportunity to broaden the advantages of text retrieval to the multi-modal scenario. Besides, we also illustrate that the language model has the ability to extract image semantics and partly map the image features to the input word embedding space. All codes are available at https://github.com/OpenMatch/MARVEL.
@@
Tianshuo Zhou1thanks: indicates equal contribution., Sen Mei [^1] footnotemark: 1, Xinze Li1, Zhenghao Liu1thanks: indicates corresponding author., Chenyan Xiong2, Zhiyuan Liu3, Yu Gu1, Ge Yu1 1Department of Computer Science and Technology, Northeastern University, China 2Language Technologies Institute, Carnegie Mellon University, United States 3Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China
@@ 1 Introduction
Multi-modal retrieval Bain et al. (2021); Awad et al. (2021); Arni et al. (2008); Chang et al. (2022) aims to return fusion results of images and texts to answer user questions. The task can be modeled using a divide-and-conquer pipeline Chang et al. (2022); Liu et al. (2023b) or universal dense retrieval Liu et al. (2023b). UniVL-DR Liu et al. (2023b) encodes queries and multi-modal documents into a universal embedding space for multi-modal retrieval. However, this work encodes image features and texts using different encoders from CLIP Radford et al. (2021) and the separated text and image encoding leads to a modality gap in representing multi-modal documents. It makes UniVL-DR design an additional image verbalization method to alleviate the modality gap and also limits the text retrieval models Karpukhin et al. (2020); Xiong et al. (2021a); Zhan et al. (2021); Li et al. (2021b); Yu et al. (2021) to excel their advantages in multi-modal scenarios.
@@ 3.1 Preliminary of Multi-Modal Retrieval
The multi-modal retrieval task requires retrievers to conduct relevance modeling, cross-modal matching, and modality fusion Liu et al. (2023b). Previous work Liu et al. (2023b) maps text and image documents in an embedding space for retrieval, encodes texts and images using different encoders, and tries to bridge the modality gap using image verbalization methods. However, this limits the capability of dense retrieval models, hindering the expansion of text matching knowledge for learning representations for multi-modal documents.
@@ 3.2 Universal Multi-Modal Encoding
We show the model architecture in Figure 2. Different from previous work Liu et al. (2023b), we can universally encode query $q$ and multi-modal documents $\mathcal{D}=\{d_{\text{Text}}^{1},...,d_{\text{Text}}^{m},d_{\text{Image}}^{1},...,d_{\text{Image}}^{n}\}$ using one encoder, T5-ANCE-CLIP:
@@ 3.2.1 Dense Retrieval with Visual Plugin
MARVEL starts from the T5-ANCE model Yu et al. (2023), which is a dense retrieval model that is well-trained using MS MARCO dataset Nguyen et al. (2016). Then we enable T5-ANCE by incorporating the visual module from the vision-language model, CLIP Radford et al. (2021), and conduct the T5-ANCE-CLIP model. We can use a universal encoder, T5-ANCE-CLIP, to encode texts, image features, and image documents.
@@ 5 Evaluation Result
In this section, we first evaluate the performance of MARVEL and conduct ablation studies. Then, we explore the effectiveness of different visual and language model fusion methods and analyze the role of visual module adaption pretraining in MARVEL. Some case studies are shown in Appendix A.8.
@@ 5.1 Overall Performance
The multi-modal retrieval performance of MARVEL and baseline models is shown in Table 2. Besides retrieval performance, we also compared retrieval efficiency in Appendix A.7.
@@ 5.2 Ablation Study
In the comparison between MARVEL and MARVEL w/o CLIP Pretraining, pretraining the visual understanding module shows its effectiveness by improving the performance on single/multi-modal retrieval tasks. It shows that the image-caption alignment relations provide some opportunities to adapt the visual module to the language model via pretraining. Subsequently, MARVEL also outperforms MARVEL w/o MS MARCO Pretraining, especially on the text retrieval task. It demonstrates that MARVEL can broaden the advantage of text relevance modeling to the multi-modal retrieval task. To unify the multi-modal encoding, MARVEL follows previous work Hannan et al. (2020) uses prompt tokens to indicate the start and end positions of encoded image features (Eq. 6), aiming to distinguish the image features from text token embeddings. These image prompt tokens bring light improvements, illustrating their roles in multi-modal document representation.
@@ 5.3 Retrieval Effectiveness of Different Visual-Language Fusion Methods
T5-CLIP (Concat) Image 37.20 81.14
@@
Table 4: Retrieval Performance of the Models using Different Visual-Language Fusion Methods. T5-CLIP (Sum/Concat) is similar to previous work Liu et al. (2023b), which only replace the image caption encoder with T5-ANCE. The CLIP-Sum model is the CLIP-DPR model from previous work Liu et al. (2023b). All models are trained with in-batch negatives. MRR@10 is used to evaluate the retrieval effectiveness of all models.
@@
Different from Liu et al. (2023b), we use T5-ANCE and CLIP as the text and image encoders, respectively. These models have different architectures and are pretrained on text retrieval and image-caption matching tasks. The multi-modal retrieval performance of CLIP-Sum decreases when we encode the image caption with a stronger retrieval model T5-ANCE (T5-CLIP-Sum) instead of CLIP. It demonstrates that incorporating an additional visual module into a well-trained dense retrieval model is still challenging for multi-modal retrieval. Notably, MARVEL provides a promising way to enable the image understanding ability of dense retrieval models by using the visual module plugin modeling method.
@@
Figure 3: Attention Distribution of MARVEL-ANCE. The attention weights of image features are shown in Figure 3(a). And the attention weight entropy of image captions and features is shown in Figure 3(b).
@@
As shown in Figure 3. The attention scores are calculated by cross attentions from the decoder to the encoder module of T5. We first show the attention weight distribution of image features in Figure 3(a). When we only finetune the language model, the attention heads tend to allocate more balanced attention weights between image features and captions, helping to adapt the visual module in the language model. On the other hand, the image features win more attention weights when the CLIP module is finetuned. However, as shown in Figure 3(b), only finetuning the CLIP module shows a scattered attention weight mechanism than other models, which misleads the T5-ANCE to capture more important information from encoded representations of documents. All these phenomena demonstrate the necessity of the training strategies of MARVEL, which pretrain visual module for adaption and only finetune the language model for multi-modal retrieval. In addition, we show the retrieval effectiveness with different finetuning methods in Appendix A.3.
Figure Text
“Hello”, “stocked”, “charged”, “institu”]
“Hi”, “auftrag”, “he”, “ban”, “sets”, “7,000”, “representation”]
“default”, “16”, “raum”, “haunt”, “unanimous”]
“russia”, “Color”, “vivid”, “HM”, “arbeit”, “default”]
5.4 Effectiveness of Visual Module Adaption Pretraining
We show some cases in Table 5 to show the effectiveness of the visual understanding module by verbalizing encoded image features. Besides, more experiments about the encoded image features are shown in Appendix A.6.
We randomly select four image documents of different topics and represent the encoded image feature with some tokens to verbalize the image semantics. Specifically, we first use the visual plugin modules of MARVEL and MARVEL w/o CLIP Pretraining to encode the image features. Then, to show the semantics of the encoded image features, we utilize cosine similarity to find the $k$-NN tokens for each encoded image feature. Finally, the tokens with the highest score are used to represent the semantics of the encoded image features.
For the first two examples, MARVEL w/o CLIP Pretraing learns more similar representations for both image documents. The related word tokens of these image documents contain lots of same tokens, such as “7,000”, “Hi”, and “RAM”, which are unrelated to the semantics of the image documents. On the contrary, our MARVEL model can learn more similar semantics to both image documents. Specifically, MARVEL verbalizes the first image document using the words “brightness”, “resident” and “store”, which are related to the image description of “Bourbon Street”. Additionally, MARVEL captures the semantics of “animals”, “wildlife” and “creatures” in the second image document. The next two instances show the effectiveness of MARVEL in learning more fine-grained semantics of the image documents by verbalizing the image documents with more related words, such as “militari”, “vehicle”, “flag”, “legislatur”, and “government”. Overall, compared with the MARVEL w/o CLIP Pretraining model, MARVEL has the ability to learn more effective representations that are closer to the semantics of the images, demonstrating the important role of MARVEL’s visual module pretraining strategy in adapting the visual understanding module for dense retrievals.
@@ 6 Conclusion
This paper proposes Multi-modAl Retrieval via Visual modulE pLugin (MARVEL). MARVEL integrates a visual plugin module with a well-trained dense retriever and pretrains the visual module with image-caption contrastive training for adaption. Our MARVEL model achieves state-of-the-art on all benchmarks by unifying the multi-modal document encoding and alleviating the modality gap between images and texts.
@@
Acknowledgments
Tianshuo Zhou, Sen Mei, Xinze Li and Zhenghao Liu are supported by the Natural Science Foundation of China under Grant (No. 62206042 and No. 62272093), the Joint Funds of Natural Science Foundation of Liaoning Province (No. 2023-MSBA-081), and the Fundamental Research Funds for the Central Universities under Grant (No. N2416012). Yu Gu is supported by the National Natural Science Foundation of China (No. U23B2019). Ge Yu is supported by the Natural Science Foundation of China under Grant (No. 62137001).
@@ Appendix A Appendix
A.1 License
We show the licenses of the datasets that we use. WebQA uses CC0-1.0 license, while ClueWeb22 shows its terms of use at website [^3] https://lemurproject.org/clueweb22/. All these licenses and agreements permit the use of their data for academic purposes.
A.2 Experimental Details of MARVEL Pretraining Data
@@
To pretrain the visual module in MARVEL, we collect the image-caption pairs from the ClueWeb22 dataset. We retain the English pages, extract the content within the image tag and use the image and alt-text to construct the image-caption pair. To ensure the quality of the pretraining dataset, following LAION-400M Schuhmann et al. (2021), we use CLIP to calculate the embeddings of images and captions and compute the cosine similarity between the two embeddings. Subsequently, we discard all samples with a cosine similarity lower than 0.3. The pretraining dataset contains 1.6M image-caption pairs, and we randomly select 10,000 pieces of data as the development set and use the rest for the pretraining visual module.
Finetune Modality MRR@10 NDCG@10 Rec@100
Text 64.89 58.71 89.98
CLIP & T5 Image 64.36 65.41 94.19
Multi 64.37 61.77 91.78
Text 64.72 58.88 90.26
T5 Image 66.12 67.49 95.12
Multi 65.15 62.95 92.40
Text 48.38 41.63 75.11
CLIP Image 56.28 56.17 87.67
Multi 49.22 45.80 80.42
@@
N/A Image 55.09 54.99 87.26
Multi 48.12 44.75 79.69
Table 6: The Retrieval Performance with Different Training Strategies. We freeze each module of MARVEL-ANCE to explore the benefits of training between different modules.
A.3 Retrieval Effectiveness of Different Finetuning Strategies
@@
As shown in Table 6, finetuning the CLIP module indeed improves the retrieval performance of the whole frozen model, especially in the image retrieval task. This observation shows that multi-modal training signals are effective to benefit the capability of visual modules. When we only tune the parameters of T5, MARVEL-ANCE achieves significant improvements over the frozen model, showing the language model’s strong ability to adapt the visual module to the dense retriever. Nevertheless, the fully finetuned model decreases the retrieval performance of MARVEL-ANCE that only finetunes T5. It shows the necessity of the pretraining-and-then-finetuning strategy of MARVEL, which pretrains the visual understanding module for adaption and finetunes the language model for multi-modal retrieval.
A.4 More Details of ClueWeb22-MM
@@
Model MRR@10 NDCG@10 Rec@100
MARVEL-DPR 55.71 52.94 88.23
w/ 1-NN Token 38.80 35.89 73.59
w/ 5-NN Tokens 42.39 39.27 75.04
w/ Random Token 37.73 35.34 71.92
MARVEL-ANCE 65.15 62.95 92.40
w/ 1-NN Token 51.37 48.27 80.47
w/ 5-NN Tokens 52.22 49.35 81.88
w/ Random Token 44.22 41.55 71.23
Setting Model Encoding Time (ms) Retrieval Time (ms)
Query Img Doc Text Doc Query
Single Modality (Text Only) BM25 - 0.1 0.1 41.9
DPR (Zero-Shot) 1.8 2.7 4.7 48.4
CLIP-Text (Zero-Shot) 0.4 0.2 0.3 32.7
Anchor-DR (Zero-Shot) 2.2 2.0 2.0 47.7
T5-ANCE (Zero-Shot) 2.1 2.0 2.0 48.3
BERT-DPR 1.1 2.7 3.1 47.6
NQ-DPR 1.3 2.8 4.9 47.8
NQ-ANCE 1.3 2.7 4.7 48.1
Divide-Conquer VinVL-DPR 1.4 5.2 4.7 48.4
CLIP-DPR 0.4 0.8 0.3 32.8
BM25 & CLIP-DPR 0.4 0.8 0.1 37.3
UnivSearch CLIP 0.4 0.8 0.3 32.8
VinVL-DPR 1.3 5.3 4.7 47.6
UniVL-DR 0.4 0.9 0.3 32.9
MARVEL 2.1 3.7 2.0 48.0
Table 12: Retrieval Efficiency. We compare the encoding and retrieval times of different architectural models on the same device. These models encode the multi-model documents and queries offline and construct the FlatL2 index using FAISS Johnson et al. (2019) for online retrieval.
To show the details of our ClueWeb22-MM dataset, we show the data collection, data processing, and data statistics in this subsection.
@@
To further guarantee the quality of the dataset, we use T5-ANCE Yu et al. (2023) to estimate the relevance between the anchor and its corresponding image document. We encode all captions of image documents using T5-ANCE, use the anchor texts as queries to retrieve the images and reserve the anchors that are ranked in the top 10. Finally, we respectively sample 10,000 queries to build the development and test set. The rest data are used for finetuning models, which contain 72,028 queries.